OpenStudy (anonymous):

h(t)=64t-16t². Compute h(1), H(2), h(4).

OpenStudy (anonymous):

That just means plug in 1, 2, and 4 for t into the expression h(t). So h(1) = 64(1) - 16(1)^2 h(2) = 64(2) - 16(2)^2 h(4) = 64(4) - 16(4)^2